If in a standard dihybrid cross (AaBb x AaBb), there were a total of 80 observed for the F2 generation, what would be the expected number for the A_B_ phenotype?
Answer #1

A punnet square for AaBb x AaBb cross geneally have 9:3:3:1 phenotype ratio however for AaBb x AaBb dihybrid cross the phenotypic ratio is 9:3:4 .

Total phenotypic ratio = 9+3+4 =16

9/16 x 80 = 45

3/16 x 80 = 15

4/16 x 80 = 20